OpenStudy (anonymous):

In a population of tree frogs, the allele that causes yellow eyes (Y) is dominant over the allele that causes orange eyes (y). If 85% of the frogs have yellow eyes, what is the frequency of the dominant allele?

OpenStudy (anonymous):

I'll assume you know the Hardy-Weinberg Principle: p is the frequency of Y in the population; q is the frequency of y in the population. p + q = 1 (1) p^2 tells us the frequency of homozygous dominants q^2 tells us the frequency of homozygous recessives 2pq tells us the frequency of heterozygotes Since 85% of the frogs have yellow eyes, we can make the following expression: p^2 + 2pq = 0.85 (2) Now its just a question of solving (1) and (2) simultaneously, which I think involves a quadratic... In any case hope this helps!

OpenStudy (anonymous):

Just realised I seriously over complicated how you work this out! If 85% have yellow eyes, then 15% must have orange eyes...and since only homozygous recessive (yy) have orange eyes, then p^2 = 0.15... so square root that and then use 1=p+q and your done, without having to solve a quadratic! Sorry for the first reply!